OpenStudy (anonymous):

QUESTION FOR THE FOLLOWING GRAPH: * Identify the coordinates of the center of this ellipse. * Use the values of a and b to locate the coordinate f the vertices. * What are the coordinates of the foci? * Graph this ellipse. Clearly label the foci and vertices

OpenStudy (anonymous):

\[\frac{ (x-2)^2 }{ 36 } + \frac{ (y+1)^2 }{ 25 } = 1\] I know that the center is (2, -1).

OpenStudy (anonymous):



OpenStudy (anonymous):

I forgot the formula for both foci and vertices. What I do know is that the foci formula below is for equations with their center at the origin. \[c^2 = a^2−b^2\] And i thought a^2 was supposed to tell me the vertices from the origin too.

OpenStudy (anonymous):

\[b ^{2}=a ^{2}\left( 1-e ^{2} \right),25=36\left( 1-e ^{2} \right)\] \[e ^{2}=1-\frac{ 25 }{36}=\frac{ 11 }{36 }\] \[e=\frac{ \sqrt{11} }{ 6 }\] \[foci are \left( \pm ae+x co-ordinate of centre,y co-ordinate of centre \right) i.e.,\left( -\sqrt{11}+2,-1 \right) and \left( \sqrt{11}+2,-11 \right)\]

OpenStudy (anonymous):

Uh oh. It went off into the distance... lolz :) But thank you!

OpenStudy (anonymous):

What is the formula for vertices when they aren't at the origin?

OpenStudy (anonymous):

i write again.they are \[\left(- \sqrt{11}+2,-1 \right) and \left( \sqrt{11}+2,-1 \right)\] \[vertices are \left( -a+x co-ordinate of centre ,y co-ordinate of centre \right) and\] \[\left( a+x co-ordinate of centre, y co-ordinate of centre \right)\]

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

FOCI: \[(−\sqrt{11}+ 2,−1)and(\sqrt{11}+ 2,−1)\] VERTICES: \[(-4,-1) and(8,-1)\]